OpenStudy (anonymous):

solve the equation 20-x/3=x/2

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

this is a bit ambiguous - do u mean (20-x)/3 = x/2?

OpenStudy (anonymous):

if so 40-2x = 3x 5x = 40 x = 8

OpenStudy (anonymous):

no, the x/3 and x/2 are supposed to be fractions but I dont know how to do them with my computer.

OpenStudy (anonymous):

ok - the problem is a lot of people forget to put bracket saround the expression so we need to multiply each term by lcd of 2 and 3 which is 6: 20-x/3=x/2 120 -2x = 3x x = 120/5 = 24

OpenStudy (anonymous):

ok?